OpenStudy (anonymous):

Which property states that the order of numbers can be changed in an addition or multiplication problem? A. Associative B. Commutative C. Transitive D. Reflexive

OpenStudy (anonymous):

B commutative eg 4 + 5 = 5 + 4 4* 5 = 5* 4

OpenStudy (anonymous):

Thank you (:

OpenStudy (anonymous):

no probs